# Deeper Diffusion Models Amplify Bias

### v1

@@ 3.2 Latent Space Abstraction Level Influences Bias-Variance Tradeoff in Diffusion Models
In this section, based on Lemma 1 and Lemma 2, we claim that the bias-variance tradeoff in diffusion models is directly influenced by the level of abstraction of the latent space representation.
@@
High abstraction levels of representations lead to “high bias”. Based on Lemma 1, highly abstracted representation makes the energy landscape corresponding to the DM’s learned memories smoother, and this smoothness results in emergence of merged minima basins. Lemma 2 establishes that converging to those merged minima basins amplifies the probability of generating majority class samples considerably. Consequently, the generated images may belong to the majority class with higher probability than the original proportion of that class. That is, the bias gets amplified.
Low abstraction levels of representations lead to “high variance.” Low abstraction levels lead to uneven energy landscapes with numerous individual local minima, which are representatives of less aggregated and more individual data points. Convergence behavior of the corrupted signal is very sensitive to the presence or absence of individual training samples in this case. As a result, this will be the state of “high variance”.
3.3 High Bias Amplifies the Inherent Bias in Datasets
@@
To verify the established relationship between high levels of abstraction and “high bias”, we designed and conducted a series of experiments to further examine this connection.
@@
We trained a series of unconditional diffusion models on custom MNIST datasets containing only digits “6” and “9” with different proportions. The base case is a diffusion model with two encoding blocks trained on a biased dataset containing 90% samples of class “6” and 10% samples of class “9”. We further trained increasingly deeper models on the same biased dataset. The models consist of three, four, five, and six encoding blocks. Interestingly, we observed that while the percentage of generated images for the base model almost aligns with the proportions of the training dataset, deeper models; which leverage increasingly more abstract representations, tend to amplify the bias. The same experiments are also conducted on datasets consisting 80-20, 70-30, and 60-40 percent of “6”s and “9”s, respectively, and the results were consistent. We also used Learned Perceptual Image Patch Similarity (LPIPS) metric [34] to measure the diversity of generated images, which showed higher diversity for deeper architectures (see Figure 4). Skeptically, this observation could be caused by over-parameterization. Hence, we also repeated the experiment on a large 2-encoding-block model that has almost the same number of parameters as the 6-encoding-block model. The results remain largely similar to the small 2-encoding-block model we experimented with earlier which confirms that it is not the number of parameters, but the abstraction level of representations that is responsible for the observed bias amplification.
@@
Figure 4: % of generated images from the minority class, and average LPIPS metric to measure their diversity. Five DMs with 2 to 6 encoding blocks each trained on four custom MNIST datasets contained on only “6”s and “9”s with biased proportions (see legend). Deeper models, that make more abstract representations, generate more diverse images (higher LPIPS) and amplify the bias. “High variance” and “High bias” states can be identified on “shallow” and “deep” models, respectively.
@@
Table 2: Percentage of images generated by unconditional DMs trained on CelebA. Higher abstraction of representations that is achieved by more number of encoding blocks $e$ leads to amplification of bias, as well as more diversity in generating images.
Dataset Proportions Number of Encoding Blocks
@@
To validate our observation further, we also trained a series of increasingly deeper DMs on CelebA dataset [20] and generated 10,000 images. As presented in Table 2, the proportions of generated images across the two classes of “Male”, “Female” support our claims. Although the proportion of “Female” samples to “Male” samples is $58.32\%$ to $41.68\%$ in the training data, the bias amplifies as the number of encoding blocks increases. Please note that the results for the classes that are combinations of genders and hair colors are not reported, because in that case the minority class (blond male) contributes to less that $1\%$ of the training data, and getting a meaningful result of its changes required generating millions of samples, so we reported the results for “gender” classes.
@@
Figure 5 shows 10 consecutive generations of two models with 4 and 6 encoding blocks, trained on CelebA starting from the same seed. In the two highlighted sample pairs, the generated images differ: the deeper model generates examples from majority classes - a “non-blond male” that makes up 40.81% of the dataset, instead of a “blond male”, which represents only 0.86% of the dataset; and a “female” that constitutes 58.32% of the dataset, instead of a “male” that accounts for 41.68%. These results are perfectly explained by our bias amplification insight.
@@
Figure 5: 10 unconditional consecutive generations of two DMs trained on CelebA starting from the same seeds: one with 6-encoding blocks and one with 4-encoding blocks. The deeper model tends to follow the biased data distribution in CelebA. Two highlighted generations are different among the two models in which the generations of the 6-encoding blocks model are in favor of the majority: non-blond male vs. blond male (40.81% vs. 0.86% ) and female vs. male (58.32% vs. 41.68%).
@@
3.4 High Variance Puts the Presumed Privacy of Diffusion Models at Risk
@@
We also investigated the state of “high variance” in which the model generates images that more closely resemble specific samples of the training data. Figure 6 shows the images generated by SD V1.5 for three prompts: “an actress”, “a queen”, and “a woman”. To force the model to work with less abstract data representations we bypassed the mid-block that contains the most abstract representation. This way, the model relies on the less abstract representations flowing via skip connections. We show the results as percentage bypassed mid-block activations in gradually increasing proportions, up to complete bypassing (100%). As can be observed, the complete bypass results in generating almost similar images for different prompt. This follows naturally from our bias-variance analysis in the previous sections. It entails that promoting higher variance of pre-trained models with bypassing, can put the (presumed) privacy of DMs at risk.
4 Training-free Generation Quality Improvement
Building further on the established bias-variance trade-off, we propose a training-free method to enhance the generation quality of existing DMs. The method exploits the insight that the images generated in “high variance” state can be more realistic because DMs can leverage memorization more easily in this state. This may come at the cost of less diversity. Nevertheless, a model can be tricked to benefit form diverse natural and realistic features during the inference process. The key to deploy this insight is in giving the model more freedom to peek into adjacent local minima during the denoising process. Hence, we propose to mask the mid-block of DMs with increasing probability along the denoising steps for improved image quality.
We use a series of random masks $\{\delta\}_{t=0}^{T}$, where $\delta_{t}$ is the mask at time step $t$, and $T$ is the total number of denoising steps with increasing sparsity such that
$\|\delta_{t}\|_{0}=\|\theta^{mid}\|_{0}\times(1-\frac{t}{T}{})^{\frac{1}{r}},\quad\text{for}\quad t=0,\dots,T,$ (11)
where .$\|.\|_{0}$ denotes pseudo-norm zero, $\delta^{mid}$ is the original parameter vector corresponding to the mid-block, and $r$ is the rate of bypassing and is set to 1 in our experiments.
By applying the mask at each time step $t$ the parameters of the mid-block is updated as follows:
$\theta_{t}^{mid}=\theta^{mid}\odot\delta_{t}.$ (12)
@@ 4 Training-free Generation Quality Improvement
Figure 6: Images generated using SD V1.5 for three closely related prompts: “an actress”, “a queen”, and “a woman”, starting from the same seed (left column). By gradually bypassing the mid-block of the model, the generations become more similar and more realistic! Bypassing the mid-block forces the model to work with less abstract representations, falling in the state of “high variance”, in which the noisy signal attends to a small number of attractors, corresponding to some specific memories.
Figure 7: (Left) The generation quality improvement method introduces additional degrees of freedom to facilitate DM models accessing more local minima with less trajectory length, leading to diversity and capturing more details from the learned memories. (Right) Image pairs generated by SDXL and the proposed method. Our method consistently yields more realistic and visually appealing outputs. Input prompt are also mentioned for each pair.
As illustrated in Figure 7 (left), bypassing weights improves degrees of freedom for the noisy signal in the energy landscape. This results in accessing the attractors more conveniently. Our method randomly changes the mask along the inference steps to avoid converging to a specific memory. Based on the established fact that in the denoising process, the overall shapes form earlier and the details emerge later [31], we start denoising without any masking, and then progressively apply increasingly larger masks. This lets the model decide on the overall shape with all its capacity which avoids distorted images, and by gradually increasing the number of zeros in the mask, we encourage the model toward a higher quality generation.
We present representative qualitative results of our method in Figure 7 (right), where SDXL’s [22] inference is modified by our method. For reference, we also include the original SDXL images using the same seed and prompt. A careful inspection reveals a consistent improvement in image perceptual quality. We observed this improvement as a general trend - see Appendix for more results. Due to the computationally subtle yet perceptually significant nature of these improvements, we directly perform human evaluation of our generated results instead of relying on debatable quantitative metrics. In Table 3, we summarize the results as percentages of the times humans preferred images generated by our method over the original SDXL counterpart using prompts from PartiPrompts dataset [33], when asked about quality and appeal of the images. In Figure 8, we also apply our technique to image-to-image generation task. We observe that our method not only maintained or improved the visual quality of the images, but it also led to more diversity in the images due to its ability to peek into multiple local minima with progressive bypassing. We provide further details and results of the experiments in the Appendix.
Table 3: Human evaluation of our method vs SDXL. The reported percentages indicate the proportion of votes preferring our method’s outputs when asked to consider visual quality and appeal. Participant were non-experts in vision, and were not provided information about the source models.
Our Method Preference Animals Objects People Places Vehicles Average
$82.22\%$$\pm 8.03$ $82.35\%$$\pm 8.32$ $78.57\%$$\pm 11.29$ $86.67\%$$\pm 9.50$ $93.33\%$$\pm 6.97$ $84.63\%$$\pm 5.65$
Figure 8: Our image enhancement applied to image-to-image generation. Our method not only maintained or improved perceptual quality, but also led to more diverse images - notice cat’s features.
@@ 5 Related Work
Explanation of Diffusion Models Many recent studies have focused on explanation of the underlying mechanisms of DMs. Yi et al. [31] showed that the generation of images in DMs start with low frequency signals in initial steps, followed by construction of high frequency signal in later steps. This behavior leads to generation of overall image first and details later in the denoising process. The memorization-generalization properties of DMs are explored in [16], [17], and [32]. Furthermore, Vastola [25] explored the elements influencing the generalization of DMs. Kwon et al. [15] uncovered the semantic latent space in DMs, which is further explored in [21] and [27]. As a more intutive alternative to common explanation of DM models’ generation as iterative denoising, Hoover et al. [11] proposed to conceptualize the inference process as memory retrieval.
Bias Mitigation in Diffusion Models DMs are accused of amplifying the biases inherent in their training data [9, 1]. However, Seshjadri et al. [24] challenged this concern and argued that in text-to-image case, bias exacerbation stems from distribution shift in training captions and prompts. Conversely, as suggested by Chen et al., the bias amplification is more serious than presumed and can potentially results in the cascading of bias towards future models[4].
Privacy of Diffusion Models To preserve the confidentiality of potentially private dataset samples, Jahanian et al. [13] proposed using generative models to have variants of samples instead of explicitly storing and working with real datasets. However, Carlini et al. [2] revealed that DMs are prone to generate images highly resembling their training data. Also it is argued that DMs can be pushed to give away their training samples using membership inference method [7]. In response to the challenge, robust DMs that address privacy concerns are proposed [6, 26, 3]. Moreover, the inherent privacy guarantees and conditions affection privacy preserving properties of datasets generated by DMs are analyzed in [29].
@@ 6 Conclusions
In this work, we explored the concept of “bias-variance trade-off” in the context of DMs, which explains bias amplification at one extreme and lack of diversity in generation at the other extreme. Our arguments are aligned with the opposing concepts of “memorization” and “generalization”, but it further captures the state beyond generalization. This study, revealed the potential disadvantages of extremely deep DMs not limited to computational cost. We also showed the vulnerability of pre-trained DMs to be pushed to “high variance” state in order to get information of their training data. The energy landscape perspective of DMs, led us to develop a training-free method to enhance the quality and details of pre-trained models.
Limitations and ethical concerns: Our contribution makes a pioneering foray into the bias-variance tradeoff in DMs. Whereas we provide concrete evidence of each insight, there are natural followup investigations that remain unexplored. For instance, optimal balance between bias-variance w.r.t. abstraction level of models, and formal treatment of ‘generalization’ on this spectrum remains unaddressed. However, our contribution lays solid groundwork for these interesting future enquiries. We also highlight that our method for enabling high variance during DM inference has the potential to be misused to reveal private training data. Extensive confirmation of this fact and remedies to preempt this, are also a part of our future endeavors.

### v2

@@ 3.2 Latent Space Abstraction Level Influences Bias-Variance Tradeoff in Diffusion Models
Based on Theorem˜1 and Theorem˜2, we claim that the bias-variance tradeoff in diffusion models is directly influenced by the level of abstraction of the latent space representation.
@@
High abstraction levels of representations lead to “high bias”. Based on Theorem˜1, highly abstracted representation makes the energy landscape corresponding to the DM’s learned memories smoother, and this smoothness results in the emergence of merged minima basins. Theorem˜2 establishes that converging to those merged minima basins amplifies the probability of generating majority class samples considerably. Consequently, the generated images may belong to the majority class with a higher probability than the original proportion of that class. That is, the bias gets amplified. Low abstraction levels of representations lead to “high variance”. Low abstraction levels lead to uneven energy landscapes with numerous individual local minima, which are representatives of less aggregated and more individual data points. The convergence behavior of the corrupted signal is very sensitive to the presence or absence of individual training samples in this case. As a result, this will be the state of “high variance”.
3.3 “High Bias State” Leads to Amplification of the Inherent Bias in Datasets
@@
To verify the established relationship between high levels of abstraction and “high bias”, we designed and conducted a series of experiments on MNIST and CelebA datasets to further examine this connection.
@@
First, we trained a series of unconditional diffusion models on custom MNIST datasets containing only digits “6” and “9” with different proportions. The base case is a diffusion model with two encoding blocks trained on a biased dataset containing 90% samples of class “6” and 10% samples of class “9”. We further trained increasingly deeper models on the same biased dataset. The models consist of three, four, five, and six encoding blocks. Interestingly, we observed that while the percentage of generated images for the base model almost aligns with the proportions of the training dataset, deeper models; which leverage increasingly more abstract representations, tend to amplify the bias. The same experiments are also conducted on datasets consisting 80-20, 70-30, and 60-40 percent of “6”s and “9”s, respectively, and the results were consistent. We also used Learned Perceptual Image Patch Similarity (LPIPS) metric Zhang et al. (2018) to measure the diversity of generated images. For each setting, we generated 1000 images and calculated the LPIPS metric for each pair. Eventually, the average LPIPS is reported, which showed higher diversity for deeper architectures (see Figure 4).
@@
Figure 4: % of generated images from the minority class, and the average LPIPS metric to measure their diversity. Five DMs with 2 to 6 encoding blocks each trained on four custom MNIST datasets contained only “6”s and “9”s with biased proportions (see legend). Deeper models, that make more abstract representations, generate more diverse images (higher average LPIPS) and amplify the bias. “High variance” and “high bias” states can be identified on “shallow” and “deep” models, respectively.
@@
Table 2: Percentage of images generated by unconditional DMs trained on CelebA. Higher abstraction of representations that is achieved by a larger number of encoding blocks leads to amplification of bias, as well as more diversity in generating images.
Dataset Proportions # of Encoding Blocks
@@
To validate our observation further, we trained a series of increasingly deeper DMs on CelebA dataset Liu et al. (2015) and generated 10,000 images. As presented in Table˜2, the proportions of generated images across the two classes of “Male”, “Female” support our claims. Although the proportion of “Female” samples to “Male” samples is $58.32\%$ to $41.68\%$ in the training data, the bias amplifies as the number of encoding blocks increases. Skeptically, this observation could be caused by over-parameterization. Hence, we repeated the experiment on a large 2-encoding-block model that has almost the same number of parameters as the 5-encoding-block model (see Table A in Appendix) . The results remain largely similar to the small 2-encoding-block model we experimented with earlier which confirms that it is not the number of parameters, but the abstraction level of representations that is responsible for the observed bias amplification. Please note that the results for the classes that are combinations of genders and hair colors are not reported, because in that case the minority class (blond male) contributes to less that $1\%$ of the training data, and getting a meaningful result of its changes required generating millions of samples, so we reported the results for “gender” classes.
@@
Figure˜5 shows 10 consecutive generations of two models with 4 and 6 encoding blocks, trained on CelebA starting from the same seed. In the two highlighted sample pairs, the generated images differ: the deeper model generates examples from majority classes - a “non-blond male” that makes up 40.81% of the dataset, instead of a “blond male”, which represents only 0.86% of the dataset; and a “female” that constitutes 58.32% of the dataset, instead of a “male” that accounts for 41.68%. These results are perfectly explained by our bias amplification insight.
@@
Figure 5: 10 unconditional consecutive generations of two DMs trained on CelebA, starting from the same seeds: one with 6-encoding blocks and one with 4-encoding blocks. The deeper model tends to follow the biased data distribution in CelebA. While most outputs are visually similar, two highlighted cases differ: the 6-encoding block model favors the majority classes, producing non-blond males over blond males (40.81 % vs. 0.86 %) and females over males (58.32 % vs. 41.68 %).
@@
3.4 “High Variance State” Undermines Privacy Preservation in Diffusion Models
Table 3: Average cosine similarity in Inception feature space and average Euclidean distance in pixel space between generated images and their corresponding $k$ nearest neighbors in the CelebA training dataset. The generated images from shallower models consistently show higher cosine similarity and lower Euclidean distance in the Inception feature and pixel spaces, respectively, indicating their greater tendency to expose training data information.
# of
Encoding
Blocks
Avg. cosSim in Inception Feature Space
for $k$ Nearest Neighbors
Avg. Euclidean Distance
in Pixel Space for $k$ Nearest Neighbors
$k=1$ $k=2$ $k=5$ $k=10$ $k=1$ $k=2$ $k=5$ $k=10$
3 0.8751 0.8612 0.8567 0.8543 97.78 99.89 113.01 116.14
4 0.8644 0.8541 0.8506 0.8420 106.15 106.25 113.11 115.03
5 0.8428 0.8355 0.8277 0.8265 106.62 106.87 107.25 117.01
6 0.8142 0.8112 0.8056 0.8016 121.41 121.84 129.91 143.05
We also investigated the state of “high variance” in which the model generates images that more closely resemble specific samples of the training data. Table 3 gives the average result of measured similarities between the generated images and their nearest neighbors in the training data. For this comparison, cosine similarity in Inception feature space and Euclidean distance in pixel space are used. The experiments are carried on for $k=1,2,5,\text{and}\ 10$. The results clearly show in general, shallower models tend to generate images resembling some specific samples of the training data more than deeper ones, as they yield larger cosine similarities between their features in Inception feature space and their nearest neighbors, while having smaller Euclidean distance in pixel space. This observation emphasizes the vulnerability of shallower models with respect to the privacy of training data in diffusion models.
Also, relevant to privacy preserving of training data is the capability of models in generating diverse images. Figure˜4 and Table 2 give average mutual LPIPS scores for generated images by models trained on MNIST and CelebA, respectively. The results consistently show lower LPIPS for shallower models, indicating their limited ability to generate diverse outputs.
@@
As a qualitative experiment, we also tried to force the pretrained model to work with less abstract data representations to simulate the “high bias” state. To do so, we bypassed the mid-block of the U-Net of SD V1.5. This way, the model relies on the less abstract representations flowing via skip connections. Figure˜6 shows the images generated by SD V1.5 for three prompts: “a woman”, “a girl”, and “an actress”. We show the results as percentage bypassed mid-block activations in gradually increasing proportions, up to complete bypassing (100%). As can be observed, the complete bypass results in generating almost similar images for different prompt. This follows naturally from our bias-variance analysis in the previous sections. It entails that promoting higher variance of pre-trained models by bypassing can put the (presumed) privacy of DMs at risk.
@@
Figure 6: Images generated with SD v1.5 for three similar prompts: “a woman”, “a girl”, and “an actress”, using the same seed (left column). Gradually bypassing the model’s mid-block makes the outputs more similar and realistic, as the model relies on less abstract representations, inducing “high-variance” state, where the noisy signal attends to only a few attractors, reflecting specific memorized patterns.
@@ 4 Related Work
Explanation of Diffusion Models Many recent studies have focused on explanation of the underlying mechanisms of DMs. Yi et al. Yi et al. (2024) showed that the generation of images in DMs starts with low frequency signals in initial steps, followed by the construction of high frequency signals in later steps. This behavior leads to the generation of the overall image first and details later in the denoising process. The memorization-generalization properties of DMs are explored in Li et al. (2023), Li et al. (2024), and Yoon et al. (2023). Furthermore, Vastola Vastola (2025) explored the elements influencing the generalization of DMs. Kwon et al. Kwon et al. (2023) uncovered the semantic latent space in DMs, which is further explored in Park et al. (2023) and Wang et al. (2025). As a more intuitive alternative to the common explanation of DM models’ generation as iterative denoising, Hoover et al. Hoover et al. (2023) proposed to conceptualize the inference process as memory retrieval.
Bias Mitigation in Diffusion Models DMs are accused of amplifying the biases inherent in their training data Friedrich et al. (2023); Bianchi et al. (2023). However, Seshjadri et al. Seshadri et al. (2023) challenged this concern and argued that in the text-to-image case, bias exacerbation stems from distribution shift in training captions and prompts. Conversely, as suggested by Chen et al., the bias amplification is more serious than presumed and can potentially results in the cascading of bias towards future modelsChen et al. (2024b).
Privacy of Diffusion Models To preserve the confidentiality of potentially private dataset samples, Jahanian et al. Jahanian et al. (2022) proposed using generative models to have variants of samples instead of explicitly storing and working with real datasets. However, Carlini et al. Carlini et al. (2023) revealed that DMs are prone to generate images highly resembling their training data. Also, it is argued that DMs can be pushed to give away their training samples using membership inference method Duan et al. (2023). In response to the challenge, robust DMs that address privacy concerns are proposed Dockhorn et al. (2023); Wang et al. (2024); Chen et al. (2024a). Moreover, the inherent privacy guarantees and conditions affecting privacy-preserving properties of datasets generated by DMs are analyzed in Wei et al. (2024).
@@ 5 Conclusions
In this work, we explored the concept of “bias-variance trade-off” in the context of DMs, which explains bias amplification at one extreme and privacy risks as well as reduced generative diversity at the other. Our arguments are aligned with the opposing concepts of “memorization” and “generalization”, but it further captures the state beyond generalization. This study revealed the potential drawbacks of extremely deep DMs, not limited to computational cost. We also showed the vulnerability of pre-trained DMs to be pushed to “high variance” state in order to get information of their training data.
Appendix A Appendix
A.1 Proof of Theorem 1
Theorem 1 (Higher abstraction in representations induces smoother energy landscapes).
$E:=E\circ(\phi^{(a)})^{-1}:Z^{(a)}\to\mathbb{R}.$ (19)
Proof.
Let $J_{\phi^{(a)^{-1}}}(z^{(a)})$ denote the Jacobian of $\phi^{(a)^{-1}}$ at $z^{(a)}$. Then, by the chain rule, the gradient of the energy function $E^{(a)}$ at $z^{(a)}$ is given by:
$\nabla E^{(a)}(z^{(a)})=J_{\phi^{(a)^{-1}}}(z^{(a)})^{\top}\nabla E(\phi^{(a)^{-1}}(z^{(a)})).$ (21)
Similarly, the Hessian of $E^{(a)}$ is given by:
$\nabla^{2}E^{(a)}(z^{(a)})=J_{\phi^{(a)^{-1}}}(z^{(a)})^{\top}\nabla^{2}E(\phi^{(a)^{-1}}(z^{(a)}))J_{\phi^{(a)^{-1}}}+\sum_{i}\frac{\partial E}{\partial x_{i}}\nabla^{2}\phi^{(a)^{-1}}(z^{(a)}).$ (22)
Taking norms on both sides, we obtain:
$\left\|\nabla^{2}E^{(a)}(z^{(a)})\right\|\leq\overbrace{\|J_{\phi^{(a)^{-1}}}\|^{2}}^{\text{decreasing in }a}.\overbrace{\left\|\nabla^{2}E\right\|}^{\text{fixed}}+\|\sum_{i}\overbrace{\frac{\partial E}{\partial x_{i}}}^{\text{fixed}}\overbrace{\nabla^{2}\phi^{(a)^{-1}}(z^{(a)})}^{\text{decreasing in }a}\|.$ (23)
Based on Lemma 1, $\phi^{(a)}$ is a contractive and smoothing map (i.e., the Jacobian norm decreases and the second derivatives of its inverse diminish with increasing $a$), both terms in the bound decrease with increasing $a$. Therefore, the norm of the Hessian satisfies
$\left\|\nabla^{2}E^{(a+1)}\right\|<\left\|\nabla^{2}E^{(a)}\right\|.$ (24)
Moreover, let $L^{(a)}$ denote the Lipschitz constant of $\nabla E^{(a)}$, so that:
$\left\|\nabla E^{(a)}(z_{1}^{(a)})-\nabla E^{(a)}(z_{2}^{(a)})\right\|\leq L^{(a)}\left\|z_{1}^{(a)}-z_{2}^{(a)}\right\|,\quad\forall z_{1}^{(a)},z_{2}^{(a)}\in Z^{(a)}.$ (25)
Since $\left\|\nabla^{2}E^{(a)}\right\|\geq L^{(a)}$, Equation 25 implies:
$L^{(a+1)}<L^{(a)}$ (26)
This shows that both the second-order and first-order variations of the energy function diminish with increasing abstraction level. Hence, the energy landscapes become progressively smoother. ∎
Lemma 1 (Higher levels of abstraction in the representation induce smaller Jacobian norm and second-order derivative of the representation mapping).
Proof.
Diminishing Jacobian: Define the sensitivity to perturbations at level $a$ as the norm of the Jacobian as:
$J(a):=\left\|\frac{\partial{z}^{(a)}}{\partial{x}}\right\|,\quad J(a+1):=\left\|\frac{\partial{z}^{(a+1)}}{\partial{x}}\right\|.$ (27)
By first-order Taylor expansion, the effect of the perturbation on the representation at level $a$ is:
$z^{(a)}(x+\delta)\approx z^{(a)}(x)+\frac{\partial{z^{(a)}}}{\partial{x}}\delta,$ (28)
and thus the magnitude of the change is:
$\left\|z^{(a)}(x+\delta)-z^{(a)}(x)\right\|\approx\left\|\frac{\partial{z^{(a)}}}{\partial{x}}\delta\right\|\leq J(a).\|\delta\|.$ (29)
Similarly, for level $a+1$:
$z^{(a+1)}(x+\delta)\approx z^{(a+1)}(x)+\frac{\partial{z^{(a+1)}}}{\partial{x}}\delta,$ (30)
and
$\left\|z^{(a+1)}(x+\delta)-z^{(a+1)}(x)\right\|\approx\left\|\frac{\partial{z^{(a+1)}}}{\partial{x}}\delta\right\|\leq J(a+1).\|\delta\|.$ (31)
As increasing abstraction decreases sensitivity to perturbations:
$\left\|z^{(a+1)}(x+\delta)-z^{(a+1)}(x)\right\|<\left\|z^{(a)}(x+\delta)-z^{(a)}(x)\right\|.$ (32)
Using the approximations in Equation˜29 and Equation˜31, this gives:
$J(a+1).\|\delta\|<J(a).\|\delta\|\implies J(a+1)<J(a).$ (33)
Diminishing second derivative of $\phi^{(a)^{-1}}$: Consider a small perturbation $\delta$ added at level $a+1$:
$z^{(a+1)}+\delta.$ (34)
The corresponding change at level $a$ is approximated by a first and second order Taylor expansion of the inverse map $\phi^{(a+1)^{-1}}$:
$z^{(a)}+\Delta z^{(a)}\approx\phi^{(a+1)^{-1}}(z^{(a+1)}+\delta)=z^{(a)}+J_{\phi^{(a+1)^{-1}}}\delta+\frac{1}{2}\delta^{\top}\nabla^{2}\phi^{(a+1)^{-1}}\delta,$ (35)
where $J_{\phi^{(a+1)^{-1}}}$ is the Jacobian matrix and $\nabla^{2}\phi^{(a+1)^{-1}}$ is the Hessian tensor of $\phi^{(a+1)^{-1}}$. The change in the energy at level $a+1$ due to $\delta$ is:
$\Delta^{2}E^{(a+1)}\approx\delta^{\top}\nabla^{2}E^{(a+1)}\delta.$ (36)
At level $a$, the change is:
$\Delta^{2}E^{(a)}\approx(J_{\phi^{(a+1)^{-1}}}\delta)^{\top}\nabla^{2}E^{(a)}(J_{\phi^{(a+1)^{-1}}}\delta),$ (37)
ignoring higher order terms involving $\nabla^{2}\phi^{(a+1)^{-1}}$. As increasing abstraction implies reduction in sensitivity, we have $|J_{\phi^{(a+1)^{-1}}}|<1$. So,
$\|\nabla^{2}E^{(a+1)}\|\leq\|J_{\phi^{(a+1)^{-1}}}\|^{2}.\|\nabla^{2}E^{(a)}\|.$ (38)
Thus,
$\|\nabla^{2}E^{(a+1)}\|<\|\nabla^{2}E^{(a)}\|$ (39)
A.2 Implementation Details
All the experiments are implemented in Pytorch and they used three NVIDIA RTX 3090 GPUs each with 24G RAM. The details of the trained models are summarized in Table˜A.
Table A: Details of trained diffusion models. Please note the number of parameters on the “large 2 encoding blocks” model, which is at the same scale of deeper models (“5 and 6 encoding blocks”). However, its bias-amplification behavior is more similar to shallower models (“2 and 3 encoding blocks”) (see Table 2 in the main paper).
Architecture
Number of
encoding blocks
Down-block types Up-block types Block out channels Layers per block # of parameters
2 DownBlock2D, AttnDownBlock2D AttnUpBlock2D, UpBlock2D 32, 64 2 1002883
2 (Large) DownBlock2D, AttnDownBlock2D AttnUpBlock2D, UpBlock2D 128, 256 2 15930883
3 DownBlock2D, AttnDownBlock2D, DownBlock2D UpBlock2D, AttnUpBlock2D, UpBlock2D 32, 64, 128 2 3692867
4
DownBlock2D, DownBlock2D, AttnDownBlock2D,
DownBlock2D
UpBlock2D, UpBlock2D, AttnUpBlock2D,
UpBlock2D
32, 32, 64, 128 2 3859203
5
DownBlock2D, DownBlock2D, DownBlock2D,
AttnDownBlock2D, DownBlock2D
UpBlock2D, AttnUpBlock2D, UpBlock2D,
UpBlock2D, UpBlock2D
32, 64, 128, 128, 256 2 16887619
6
DownBlock2D, DownBlock2D, DownBlock2D,
DownBlock2D, AttnDownBlock2D, DownBlock2D
UpBlock2D, AttnUpBlock2D, UpBlock2D,
UpBlock2D, UpBlock2D, UpBlock2D
32, 64, 128, 128, 256, 256 2 27290691
A.3 Bias Amplification
Figure˜A and Figure˜B show qualitative examples of bias amplification in deeper models. Figure˜A shows representative examples where shallower models generated images of “blond-male” minority group, but the images got changed to the majority groups (classes) of “non-blond male” and “blond female” for the deeper models when starting from the same seeds. Similarly, Figure˜B presents cases in which shallow models generate “male” images that are minority group, while deeper models converted the same seeds to the majority group, i.e., “female”.
Figure A: Bias amplification in deeper models. Top row shows samples of generated images resembling “blond male” by a relatively shallow model consisting of three encoding blocks. Lower rows show generated images of increasingly deeper models starting from the same seeds. It can be seen that the deeper models prefer generating samples from the majority groups (non-blond male and blond-female) instead of the minority group (blond-male), thereby amplifying the data bias.
Figure B: Representative cases where shallow models generate samples of minority group “male”, but deeper models generate majority group “female” starting from the same seeds.
A.4 Pushing Models to High Variance State
Illustrated in Figure˜C, the model generates slightly different images for prompts “an actress”, “a queen”, , and “a woman”. Then, by gradually bypassing the mid-block, the model gets pushed to “high variance” state. Correspondingly, the generated images begin to look more alike and realistic in the “high variance” state, supporting our claim of attending to lower number of memories in this state.
Figure C: A model pushed towards “high variance” state by bypassing its mid-block generates similar looking images for slightly different prompts. The resulting images not only have high perceptual similarity across the prompts, but also better realism. This behavior is consistent with our arguments about “high variance” state, in which the model attends to fewer memories.